In a system of two crossed polarisers, it is found that the intensity of light from the second
polarizer is half from that of first polarizer. The angle between their pass axes is
a)pi/4 b)pi/6 c)pi/3 d)pi/2

Grade:12

1 Answers

Vikas TU
6 years ago
As we all know that
I(final) = I (initial) cos2Ѳ
I /2= I cos2Ѳ
½ = cos2Ѳ
cos Ѳ = 1/√2
Ѳ = pi/4
Option 1 is correct